OpenStudy (anonymous):

1+5(x-8) is greater than or equal to 2-(x+5)

OpenStudy (anonymous):

What are possible choices if any?

OpenStudy (mayankdevnani):

1 sol.=1+5x-40=5x-39 2 sol.=2-x-5=-x-3

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):